OpenStudy (anonymous):

un has 1200 feet of fencing materials to build a pen for his sheep and goats. The pen will be a rectangle with a divider down the middle. What should be the maximum are for the pen?

hartnn (hartnn):

the fence perimeter = 1200 let the length be 'x' and breadth be 'y' so, 2(x+y) =1200 y= 600-x now Area =A=xy =x(600-x) to get max. area, differentiate A and set it =0

hartnn (hartnn):

A = 600x -x^2 since you can take simple derivates, can you derivate A ?

OpenStudy (anonymous):

600-2x

OpenStudy (anonymous):

is it right?

hartnn (hartnn):

now set that =0 and find x

hartnn (hartnn):

yes, thats correct ! good :)

hartnn (hartnn):

now set that =0 and find x

OpenStudy (anonymous):

300

hartnn (hartnn):

yes, now Area = x(600-x) =... ?

OpenStudy (anonymous):

90000

hartnn (hartnn):

\(\checkmark \)

OpenStudy (anonymous):

is that the answer already?

hartnn (hartnn):

yes, you needed max area...dimensions we found by differentiating, and then we plugged it into area and got 90000 sq. ft.